# Speedups of linearly recurrent subshifts

### v1

@@ 2 Preliminaries and miscellaneous results
Using the above notation, linear recurrence means that for every $w\in{\mathcal{L}}(X)$ and $x\in X$ and $i\in{\mathbb{Z}}$, there is $k\in\{0,1,\dots,L|w|\}$ such that $x_{i+k+1}\dots x_{i+k+|w|}=w$. Linear recurrence is preserved by topological conjugacy, although the constant $L$ can change. Linearly recurrent shifts are automatically minimal and in fact uniquely ergodic. Let $\mu_{X}$ denote the corresponding shift-invariant probability measure.
2.1 Subshifts of finite type and sofic shifts
@@ 2.2 Word-complexity of speedups
Given a homeomorphic speedup $S=\sigma^{p}$ of a Cantor system $(X,\sigma)$, there is a constant $K$ such that the word-complexity of $S$ satisfies $\mathfrak{p}_{S}(n)\leq K\,\mathfrak{p}_{\sigma}(p_{\max}\,n)$ for $p_{\max}=\max\{p(x):x\in X\}$ from (1).
@@ 3 Finite non-abelian group extensions
$G_{x}:=\{g\in G:\exists n_{k}\nearrow\infty\text{ such that }\phi^{n_{k}}(x)\equiv g\text{ and }T^{n_{k}}(x)\to x\}.$
@@
However, this still doesn’t imply that $G_{x}$ is abelian, because $(n_{k}+n^{\prime}_{l})_{l\in{\mathbb{N}},k\in K}$ and $(n^{\prime}_{l}+n_{k})_{k\in{\mathbb{N}},l\in L}$ might be entirely different sequences for finite set $K,L\subset{\mathbb{N}}$, with only finite overlap. Letting $k\to\infty$ and $l\to\infty$ at different speeds can still lead to different limit points in $G$. It seems that if $T$ is equicontinuous, then $G_{x}$ is an abelian subgroup.
@@
Let $x,y\in X$ be two points in the same $T$-orbit. Then $G_{x}$ and $G_{y}$ are conjugate subgroups of $G$, i.e., there is $g$ such that $G_{y}=g\cdot G_{y}\cdot g^{-1}$. We can in fact take $g=\phi^{n}(y)$ where $n\in{\mathbb{Z}}$ is such that $T^{n}(y)=x$.
@@
The map $x\mapsto G_{x}$ is piecewise constant.
Proof. Let $x\in X$ be “minimal” in the sense that $G_{y}$ is not a proper subgroup of $H:=G_{x}$ for every $y\in X$. For each $g\in G$, we can try to find a sequence $(r^{g}_{k})_{k\in{\mathbb{N}}}$ such that $\phi^{r^{g}_{k}}(x)\equiv g$ and $d(T^{r^{g}_{k+1}}(x),x)\leq\frac{1}{2}d(T^{r^{g}_{k}}(x),x)$ for all $k$. If $g\notin H$, then such a sequence is necessarily finite, so there is a largest integer $N_{g}$ in this sequence. (If $\phi^{n}(x)\neq g$ for all $n\geq 0$, then we set $N_{g}=0$.)
@@
By minimality, there is $N\in{\mathbb{N}}$ such that $X=\bigcup_{i=0}^{N-1}$, and by Lemma 3.1, $G_{z}=g\cdot H\cdot g^{-1}$ for each $z\in T^{i}(U_{x}\cap A_{i})$, $i=0,\dots,N-1$, where $A_{i}$ is any set on which $\phi^{i}\equiv g$ is constant. This proves that $x\mapsto G_{x}$ is piecewise constant. $\Box$
@@
$\mu\left(A\cap T^{n}A\cap\{x\in X:\phi^{n}(g,x)=h\}\cap\{x\in X:T^{n}(x)\neq x\})\right)>0.$
For abelian groups $G$, $G_{x}$ coincides with the set of essential values, but for non-abelian groups, the essential values may not be more than the centralizer of $G$, and thus of less use.
@@
Assume that the $T$-invariant measure $\mu$ is ergodic. If the subgroup of essential values $H=G_{x}$ for all $x\in X$, then $\mu\otimes\nu_{H}$ is invariant and ergodic for the skew-product $F:X\times H\to X\times H$, where $\nu_{H}$ is right Haar measure on $H$.
We give the proof for finite groups, so right Haar measure $\nu_{H}$ is actually counting measure on $H$. In this case, $\mu_{X}\otimes\nu_{H}$ is $F$-invariant.
Proof. Note that the skew-product preserves $\mu\otimes\nu_{H}$. To prove ergodicity of this measure, assume that $Y\subset X\times H$ is invariant under the skew-product and $\mu\otimes\nu_{H}(Y)>0$. Since $\mu$ is ergodic, the projection of $Y$ onto $X$ has full $\mu$-measure.
Assume now by contradiction that $\mu\otimes\nu_{H}(Y^{c})>0$. Both $Y$ and $Y^{c}$ are invariant, so their projections are $T$-invariant sets of $X$ of positive measure, and by ergodicity of $\mu$, these projections are $X\bmod\mu$. Therefore there is a set $A\subset X$ and $h_{1},h_{2}\in H$ such that
@@
Take $h=h_{1}^{-1}\cdot h_{2}\in H$, so by definition of essential value, there is $n\in{\mathbb{Z}}$ such that $\mu(B)>0$ for $B:=A\cap T^{n}(A)\cap\{x\in X:\phi^{n}(x)=h,x\neq T_{g}(x)\}$. But then
@@
contradicting the choice of $A$ and $h_{2}$ in (4). This argument also shows that $\operatorname{orb}_{F}(x,h)$ is dense in $X\times H$ for $\mu_{X}$-a.e. $X$ and every $h\in H$, finishing the proof. $\Box$
@@
Let $(X,T,\mu)$ be uniquely ergodic and $G$ a compact group with (right) Haar measure $\nu_{G}$. Let the group extension $S:Y\to Y$ defined on $Y:=X\times G$ as $F(x,g)=(T(x),g\phi(x))$ for some $\phi:X\to G$. If $F$ is ergodic w.r.t. $\nu=\mu_{X}\otimes\nu_{G}$, then $F$ is uniquely ergodic.
Proof. Let $(x,g)\in Y$ be the $\mu_{X}\otimes\nu_{G}$-typical point, so it satisfies Birkhoff’s Ergodic Theorem w.r.t. every continuous function $\varphi:Y\to{\mathbb{R}}$. For any $h\in G$, $\varphi_{h}$ defined by $\varphi_{h}(g)=\varphi(hg)$ is continuous too, so $(x,hg)$ is $\nu$-typical w.r.t. $\varphi$ because $(x,g)$ is $\nu$-typical w.r.t. $\varphi_{h}$. It follows that there is a subset $W\subset X$ with $\mu(W)=1$ such that $W\times G$ consists entirely of $\nu$-typical points.
@@ 4 Linear recurrence of speedups
where ${\mathcal{R}}:=\dots R_{e_{-2}}R_{e_{-1}}R_{e_{0}}R_{e_{1}}R_{e_{2}}\dots\in X_{w}$ is the concatenation of return words representing $x\in X$. In fact, $(X_{R},\sigma)$, the so-called derived shift of $(X,\sigma)$, is a linearly recurrent shift by itself. Despite there being infinitely many words $w$, there are only finitely many different derived shifts up to conjugacy, see [8].
@@
An outline of the proof of why linear recurrence of $(X,\sigma)$ implies linear recurrence of the speedup starts as follows. Each bi-infinite sequence $x\in X$ can be written as concatenation ${\mathcal{R}}=\dots R_{-2}R_{-1}R_{0}R_{1}R_{2}\dots$ of return words associated to some subword $w$ of $x$. We take $w$ so long that the return words contain the essential information of the $S$-orbit that pass through it. That is, the $c$ different $S$-orbits that $\operatorname{orb}(x)$ splits into, divides the entries of $R_{i}$ into $c$ subsets, characterized by so-called entry positions. For each pair of concatenated return words $R_{i}R_{i+1}$, the transit of the $j$-th $S$-orbit, starting at the $j$-th entry position in $R_{i}$, uniquely selects an entry position in $R_{i+1}$, say $j^{\prime}$. Doing this for all $j\in\{1,\dots,c\}$, we get a permutation $\psi_{R_{i}R_{i+1}}:\{1,\dots,c\}\to\{1,\dots,c\}$. This leads to a group extension, i.e, skew-product of $X_{w}\times{\mathfrak{S}}$, where $X_{w}$ is the shift space expressed in return words of $w$, and $\mathfrak{S}$} is the group of permutations of $\mathfrak{S}$}. This is the first step of the proof, which we formalize in the following construction.
@@
by setting $\psi_{RR^{\prime}}(s)=j^{\prime}$ if the $S$-orbit passing through the $j$-th entry position in the prefix word $w$ of $Rw$ also passes through the $j^{\prime}$-th entry position of the prefix word $w$ in $R^{\prime}$. Let $\mathfrak{S}$} denote the collection of permutations of $\{1,\dots,c\}$. Thus, if the $c$ orbits (numbered $1,\dots,c$) enter the prefix word $w$ of the “first” letter $R_{e_{0}}$ of $\mathcal{R}$} from the left to the right (which is indicated as the indentity permutation $e\in{\mathfrak{S}}$), then they enter the prefix word $w$ of $R_{e_{1}}$ as $\psi_{R_{e_{0}}R_{e_{1}}}$, and they enter the prefix word $w$ of $R_{e_{1}}$ as $\psi_{R_{e_{0}}R_{e_{1}}}\psi_{R_{e_{1}}R_{e_{2}}}$, etc. Thus the entries of the distinct orbits is described by a skew-product
$F:X_{w}\times{\mathfrak{S}}\to X_{w}\times{\mathfrak{S}},\qquad({\mathcal{R}},s)\mapsto(\sigma({\mathcal{R}}),\ s\cdot\psi_{R_{0}R_{1}}).$
@@
There is $L^{*}$ depending only on $L$ such that for every $w\in{\mathcal{L}}(X)$ and ${\mathcal{R}}\in X_{w}$, every $s\in{\mathfrak{S}}$ reoccurs within $L^{*}|w|$ iterates of $F$.
Before proving Lemma 4.1 formally, we give an idea of the proof. First, we study the shapes of the transitive components of $X_{w}\times{\mathfrak{S}}$. To find such component, choose $(x_{0},s_{0})\in X_{w}\times{\mathfrak{S}}$ and take its orbit closure, say $X_{0}$. By minimality of $(X_{w},\sigma)$, $X_{0}$ projects to the whole of $X_{w}$ in the base direction, but in the fiber direction, for each $x\in X_{w}$, $X_{0}$ projects to $\{x\}\times{\mathfrak{S}}_{x}$, where ${\mathfrak{S}}_{x}$ is a subgroup of $\mathfrak{S}$}. The assignment $x\mapsto{\mathfrak{S}}_{x}$ is continuous (as shown in Proposition 3.1), and in fact piecewise constant because $\mathfrak{S}$} is discrete. Hence we can take a (cylinder) subset $Z\subset X_{w}$, such that ${\mathfrak{S}}^{\prime}:={\mathfrak{S}}_{x}$ is constant for all $x\in Z$. This means that the return map to $Z$ is in fact the derive shift $(X_{w^{\prime}},\sigma)$.
Replacing the shift $\sigma$ on $X_{w}$ by the first return map to $Z$, we can assure that the skew-product on $(X_{w^{\prime}},{\mathfrak{S}}^{\prime})$ is transitive. Now Proposition 3.2 implies that this skew-product is also minimal, so for each $s\in{\mathfrak{S}}^{\prime}$, the set $([w^{\prime}],s)$ reoccurs with a finite gap.
@@
Proof of Lemma 4.1. From Proposition 3.1 we know that $F$ preserves the transitive subset $\{(x,{\mathfrak{S}}_{x}):x\in X_{\omega}\}$, where ${\mathfrak{S}}_{x}$ are (conjugate) subgroups of $\mathfrak{S}$} and $x\mapsto{\mathfrak{S}}_{x}$ is piecewise continuous, hence piecewise constant. Therefore we can take a longer word $w^{\prime}$ (associated with a smaller cylinder set $Z\subset X_{w}$) on which ${\mathfrak{S}}_{x}\equiv{\mathfrak{S}}^{\prime}$ is a constant subgroup of $\mathfrak{S}$}. Without loss of generality, we can assume that $Z$ is given by a concatenation of return words of $w$ that itself is return word of some word $w^{\prime}$ in the original shift space $X$ (naturally, $w^{\prime}$ is longer than $w$). Thus the return map to $Z$ is the same as $(X_{w^{\prime}},\sigma)$ and the group-extension $F:X_{w^{\prime}}\times{\mathfrak{S}}^{\prime}\to X_{w^{\prime}}\times{\mathfrak{S}}^{\prime}$ is transitive.
@@
where $w_{j}=x_{p_{j}(x)}x_{p_{j}(x)}\dots w_{p_{j+1}(x)-1}$ and $n_{1},n_{2}\leq N$ are chosen minimal such that $p_{-n_{1}}(x)\leq-N$ and $p_{k+n_{2}}(x)\geq|v|+N$. The word $\bar{w}=w_{-n_{1}}\dots w_{-1}w_{0}w_{1}\dots w_{k}w_{k+1}\dots w_{k+n_{1}}$ reappears in every $S$-orbit with gap $L|w|\leq Lp_{\max}(|v|+2N)\leq L^{\prime}|v|$ for some appropriate $L^{\prime}$ independent of $v$. But whenever $w$ appears in an $S$-orbit, $v$ appears as subword in the corresponding reconstructed $\sigma$-orbit. This shows that $(X,\sigma)$ is linearly recurrent with constant $L^{\prime}$. $\Box$
Acknowledgement HB was supported by the ANR-FWF Project I 6750-N. He also thanks Klaus Schmidt for useful discussions on non-abelian group extensions.

### v2

@@ 2.2 Word-complexity of speedups
In the above notation of subshifts, linear recurrence means that for every $w\in{\mathcal{L}}(X)$ and $x\in X$ and $i\in{\mathbb{Z}}$, there is $k\in\{0,1,\dots,L|w|\}$ such that $x_{i+k+1}\dots x_{i+k+|w|}=w$. Linear recurrence is preserved by topological conjugacy, although the constant $L$ can change. Linearly recurrent shifts are automatically minimal and in fact uniquely ergodic.
@@
Given a homeomorphic speedup $S=\sigma^{p}$ of a subshift $(X,\sigma)$, there is a constant $K$ such that the word-complexity of $S$ satisfies $\mathfrak{p}_{S}(n)\leq K\,\mathfrak{p}_{\sigma}(p_{\max}\,n)$ for $p_{\max}=\max\{p(x):x\in X\}$ from (1).
@@ 3 Finite non-abelian group extensions
$G_{x}:=\{g\in G:\exists n_{k}\nearrow\infty\text{ such that }\phi^{n_{k}}(x)\equiv g\text{ and }T^{n_{k}}(x)\to x\}.$ (3)
@@
However, this still doesn’t imply that $G_{x}$ is abelian, because $(n_{k}+n^{\prime}_{l})_{l\in{\mathbb{N}},k\in K}$ and $(n^{\prime}_{l}+n_{k})_{k\in{\mathbb{N}},l\in L}$ might be entirely different sequences for infinite sets $K,L\subset{\mathbb{N}}$, with only finite overlap. Letting $k\to\infty$ and $l\to\infty$ at different speeds can still lead to different limit points in $G$. It seems that if $T$ is equicontinuous, then $G_{x}$ is an abelian subgroup.
@@
Let $x,y\in X$ be two points in the same $T$-orbit. Then $G_{x}$ and $G_{y}$ are conjugate subgroups of $G$, i.e., there is $g$ such that $G_{y}=g\cdot G_{x}\cdot g^{-1}$. We can in fact take $g=\phi^{n}(y)$ where $n\in{\mathbb{Z}}$ is such that $T^{n}(y)=x$.
@@
Assume that $T:X\to X$ is minimal. The map $x\mapsto G_{x}$ is piecewise constant.
Proof. Let $x\in X$ be “minimal” in the sense that for no $y\in X$, $G_{y}$ is a proper subgroup of $H:=G_{x}$. For each $g\in G$, we can try to find a sequence $(r^{g}_{k})_{k\in{\mathbb{N}}}$ such that $\phi^{r^{g}_{k}}(x)\equiv g$ and $d(T^{r^{g}_{k+1}}(x),x)\leq\frac{1}{2}d(T^{r^{g}_{k}}(x),x)$ for all $k$. If $g\notin H$, then such a sequence is necessarily finite, so there is a largest integer $N_{g}$ in this sequence. (If $\phi^{n}(x)\neq g$ for all $n\geq 0$, then we set $N_{g}=0$.)
@@
By minimality, there is $\ell\in{\mathbb{N}}$ such that $X=\bigcup_{i=0}^{\ell-1}U_{x}$, and by Lemma 3.1, $G_{z}=g\cdot H\cdot g^{-1}$ for each $z\in T^{i}(U_{x}\cap A_{i})$, $i=0,\dots,\ell-1$, where $A_{i}$ is any set on which $\phi^{i}\equiv g$ is constant. This proves that $x\mapsto G_{x}$ is piecewise constant. $\Box$
@@
$\mu\left(A\cap T^{-n}A\cap\{x\in X:\phi^{n}(x)=h\}\cap\{x\in X:T^{n}(x)\neq x\})\right)>0.$
For abelian groups $G$, $G_{x}$ coincides with the set of essential values, but for non-abelian groups, the essential values may not be more than the centralizer of $G$, or the intersection $\bigcap_{x\in X}G_{x}$, and thus $G$ is of less use.
@@
Let $T:X\to X$ be minimal with an ergodic $T$-invariant measure $\mu$. Let $F(x,g)=(T(x),g\cdot\phi(x))$ be a continuous skew-product, for a finite group $G$. Then there is $x\in X$, a neighbourhood $U$ of $x$ and an ergodic $F$-invariant measure $\rho$ such that $\rho|_{U\times H}$ is a multiple of $\mu|_{U}\otimes\nu_{H}$, where $H=G_{x}$ and $\nu_{H}$ is right Haar measure (i.e., counting measure) on $H$.
Proof. The proof of Proposition 3.1 implies that there is $x\in X$ and a neighbourhood $U=U_{x}$ of $x$ such that if $T^{n}(y)\in U$ for some $y\in U$, then $\phi^{n}(y)\in H$. Let $T_{U}=T^{\tau}$ be the first return map to $U$, where $\tau(y)=\min\{n\geq 1:T^{n}(y)\in U\}$ is the first return time of $y\in U$ to $U$. Then $\mu_{U}:=\frac{1}{\mu(U)}\mu|_{U}$ is $T_{U}$-invariant and ergodic because $\mu$ is $T$-invariant and ergodic. Let
$F_{U}:U\times H\to U\times H,\qquad(y\,,\,g)\mapsto(T_{U}(y)\ ,\ g\cdot\phi^{\tau(y)}(y))$
be the corresponding skew-product; it preserves $\mu_{U}\otimes\nu_{H}$.
To show ergodicity, assume that $Y\subset U\times H$ is invariant under the skew-product and $\mu_{U}\otimes\nu_{H}(Y)>0$. Since $\mu_{U}$ is ergodic, the projection of $Y$ onto $U$ has full $\mu_{U}$-measure.
Assume now by contradiction that $\mu_{U}\otimes\nu_{H}(Y^{c})>0$. Both $Y$ and $Y^{c}=(U\times H)\setminus Y$ are invariant, so their projections are $T_{U}$-invariant sets of $U$ of positive measure, and by ergodicity of $\mu$, these projections are $U\bmod\mu_{U}$. Therefore there is a set $A\subset U$ and $h_{1},h_{2}\in H$ such that
@@
(We could write $T$ here instead of $T_{U}$ because $T^{n}(y)\in U$ for $y\in U$ is equivalent to $T^{n}(y)=T^{n^{\prime}}_{U}(y)$ for some $|n^{\prime}|\leq|n|$.) Take $h=h_{1}^{-1}\cdot h_{2}\in H$, so by definition of essential value, there is $n\in{\mathbb{Z}}$ such that $\mu(B)>0$ for $B:=A\cap T^{-n}(A)\cap\{x\in X:\phi^{n}(x)=h,x\neq T^{n}(x)\}$. But then
@@
contradicting the choice of $A$ and $h_{2}$ in (5). Hence $\mu_{U}\otimes\nu_{H}$ is ergodic.
Finally, using Kac’ formula, we spread this mass over the whole skew-product space to obtain the required $F$-invariant measure $\rho$:
$\rho(\,\cdot\,):=\frac{1}{\Lambda}\sum_{n\geq 1}\mu_{U}\otimes\nu_{H}\left(F^{-n}(\,\cdot\,)\cap\{(y,g)\in U\times H:n<\tau(y)\}\right),$
for normalising constant $\Lambda:=\sum_{n\geq 1}n\mu_{U}(\{y:\tau(y)=n\})$. This formula gives $\rho|_{U\times H}=\frac{1}{\Lambda}\mu_{U}\otimes\nu_{H}$, and the ergodicity of $\rho$ follows from the ergodicity of $\mu_{U}\otimes\nu_{H}$. $\Box$
@@
Let $(X,T,\mu)$ be uniquely ergodic and $G$ a compact group with (right) Haar measure $\nu_{G}$. Let the group extension $S:Y\to Y$ defined on $Y:=X\times G$ as $F(x,g)=(T(x),g\phi(x))$ for some $\phi:X\to G$. If $F$ is ergodic w.r.t. $\nu=\mu\otimes\nu_{G}$, then $F$ is uniquely ergodic.
Proof. Let $(x,g)\in Y$ be the $\mu\otimes\nu_{G}$-typical point, so it satisfies Birkhoff’s Ergodic Theorem w.r.t. every continuous function $\varphi:Y\to{\mathbb{R}}$. For any $h\in G$, $\varphi_{h}$ defined by $\varphi_{h}(g)=\varphi(hg)$ is continuous too, so $(x,hg)$ is $\nu$-typical w.r.t. $\varphi$ because $(x,g)$ is $\nu$-typical w.r.t. $\varphi_{h}$. It follows that there is a subset $W\subset X$ with $\mu(W)=1$ such that $W\times G$ consists entirely of $\nu$-typical points.
@@ 4 Linear recurrence of speedups
where ${\mathcal{R}}:=\dots R_{e_{-2}}R_{e_{-1}}R_{e_{0}}R_{e_{1}}R_{e_{2}}\dots\in X_{w}$ is the concatenation of return words representing $x\in X$. In fact, $(X_{w},\sigma)$, the so-called derived shift of $(X,\sigma)$, is a linearly recurrent shift by itself. Despite there being infinitely many words $w$, there are only finitely many different derived shifts up to conjugacy, see [8].
@@
An outline of the proof of why linear recurrence of $(X,\sigma)$ implies linear recurrence of the speedup starts as follows. Each bi-infinite sequence $x\in X$ can be written as concatenation ${\mathcal{R}}=\dots R_{e_{-2}}R_{e_{-1}}R_{e_{0}}R_{e_{1}}R_{e_{2}}\dots$ of return words associated to some subword $w$ of $x$. We take $w$ so long that the return words contain the essential information of the $S$-orbit that pass through it. That is, the $c$ different $S$-orbits that $\operatorname{orb}(x)$ splits into, divides the entries of $R_{i}$ into $c$ subsets, characterized by so-called entry positions. For each pair of concatenated return words $R_{e_{i}}R_{e_{i+1}}$, the transit of the $j$-th $S$-orbit, starting at the $j$-th entry position in $R_{e_{i}}$, uniquely selects an entry position in $R_{e_{i+1}}$, say $j^{\prime}$. Doing this for all $j\in\{1,\dots,c\}$, we get a permutation $\psi_{R_{e_{i}}R_{e_{i+1}}}:\{1,\dots,c\}\to\{1,\dots,c\}$. This leads to a group extension, i.e, skew-product of $X_{w}\times{\mathfrak{S}}$, where $X_{w}$ is the shift space expressed in return words of $w$, and $\mathfrak{S}$} is the group of permutations of $\mathfrak{S}$}. This is the first step of the proof, which we formalize in the following construction.
@@
by setting $\psi_{RR^{\prime}}(s)=j^{\prime}$ if the $S$-orbit passing through the $j$-th entry position in the prefix word $w$ of $Rw$ also passes through the $j^{\prime}$-th entry position of the prefix word $w$ in $R^{\prime}w$. Let $\mathfrak{S}$} denote the collection of permutations of $\{1,\dots,c\}$. Thus, if the $c$ orbits (numbered $1,\dots,c$) enter the prefix word $w$ of the “first” letter $R_{e_{0}}w$ of $\mathcal{R}$} from the left to the right (which is indicated as the indentity permutation $e\in{\mathfrak{S}}$), then they enter the prefix word $w$ of $R_{e_{1}}w$ as $\psi_{R_{e_{0}}R_{e_{1}}}$, and they enter the prefix word $w$ of $R_{e_{1}}w$ as $\psi_{R_{e_{0}}R_{e_{1}}}\psi_{R_{e_{1}}R_{e_{2}}}$, etc. Thus the entries of the distinct orbits is described by a skew-product
$F:X_{w}\times{\mathfrak{S}}\to X_{w}\times{\mathfrak{S}},\qquad({\mathcal{R}},s)\mapsto(\sigma({\mathcal{R}}),\ s\cdot\psi_{R_{e_{0}}R_{e_{1}}}).$
@@
There is $L^{*}$ depending only on $L$ such that for every $w\in{\mathcal{L}}(X)$ and ${\mathcal{R}}\in X_{w}$, every $s\in{\mathfrak{S}}$ reoccurs within $L^{*}|w|$ iterates of $F$, i.e., there is $1\leq n\leq L^{*}|w|$ such that $F^{n}({\mathcal{R}},s)=(\sigma^{n}({\mathcal{R}}),s)$.
Before proving Lemma 4.1 formally, we give an outline of the proof. First, we study the shapes of the transitive components of $X_{w}\times{\mathfrak{S}}$. To find such component, choose $(\hat{x},\hat{s})\in X_{w}\times{\mathfrak{S}}$ and take its orbit closure, say $Y$. By minimality of $(X_{w},\sigma)$, $Y$ projects to the whole of $X_{w}$ in the base direction, but in the fiber direction, for each $x\in X_{w}$, $Y$ projects to $\{x\}\times s_{x}{\mathfrak{S}}_{x}$, where ${\mathfrak{S}}_{x}$ is a subgroup of $\mathfrak{S}$}, the analog of $G_{x}$ in (3) and $s_{x}\in{\mathfrak{S}}$. The assignment $x\mapsto{\mathfrak{S}}_{x}$ is continuous, and piecewise constant because $\mathfrak{S}$} is discrete. The proof of Proposition 3.1, in fact, gives a (cylinder) subset $U\subset X_{w}$, such that both $s_{x}\equiv e$ and ${\mathfrak{S}}^{\prime}:={\mathfrak{S}}_{x}$ is constant for all $x\in U$. This means that the first return map to $U$ is in fact the derived shift $(X_{w^{\prime}},\sigma)$.
Replacing the shift $\sigma$ on $X_{w}$ by the first return map to $U$, we can assure that the skew-product on $(X_{w^{\prime}},{\mathfrak{S}}^{\prime})$ is transitive and ergodic by Lemma 3.2. Now Proposition 3.2 implies that this skew-product is also minimal, so for each $s\in{\mathfrak{S}}^{\prime}$, the set $([w^{\prime}],s)$ reoccurs with a finite gap.
@@
Proof of Lemma 4.1. Let $Y=\bigcup_{x\in X_{w}}\{x\}\times s_{x}{\mathfrak{S}}_{x}$ be a transitive $F$-invariant subset of $X_{w}\times{\mathfrak{S}}$. From Proposition 3.1, we can take a word $w^{\prime}$ (longer than $w$ and associated with a cylinder set $U\subset X_{w}$) on which $s_{x}\equiv e$ and ${\mathfrak{S}}_{x}\equiv{\mathfrak{S}}^{\prime}$ is a constant subgroup of $\mathfrak{S}$}. Without loss of generality, we can assume that $U$ is given by a concatenation of return words of $w$ that itself is return word of some word $w^{\prime}$ in the original shift space $X$. Thus the return map to $U$ is the same as $(X_{w^{\prime}},\sigma)$ and the group-extension $F:X_{w^{\prime}}\times{\mathfrak{S}}^{\prime}\to X_{w^{\prime}}\times{\mathfrak{S}}^{\prime}$ is transitive.
@@
where $w_{j}=x_{p_{j}(x)}x_{p_{j}(x)+1}\dots x_{p_{j+1}(x)-1}$ and $n_{1},n_{2}\leq N$ are chosen minimal such that $p_{-n_{1}}(x)\leq-N$ and $p_{k+n_{2}}(x)\geq|v|+N$. The word $\bar{w}=w_{-n_{1}}\dots w_{-1}w_{0}w_{1}\dots w_{k}w_{k+1}\dots w_{k+n_{1}}$ reappears in every $S$-orbit with gap $L|w|\leq Lp_{\max}(|v|+2N)\leq L^{\prime}|v|$ for some appropriate $L^{\prime}$ independent of $v$. But whenever $w$ appears in an $S$-orbit, $v$ appears as subword in the corresponding reconstructed $\sigma$-orbit. This shows that $(X,\sigma)$ is linearly recurrent with constant $L^{\prime}$. $\Box$
Acknowledgements HB was supported by the ANR-FWF Project I 6750-N. He also thanks Klaus Schmidt for useful discussions on non-abelian group extensions, and Jiao Changhua for communicating an issue with an earlier version of Lemma 3.2.